# A Fast Algorithms for Fourier extension based on boundary interval data

### v1

@@ 4.2 Testing the parameter $m_{\Delta}$
The size of parameter $m_{\Delta}$ determines whether this study is meaningful. This subsection tests the influences of other parameters on it. First, we take $M>{T}_{\Delta}\gamma\omega$ for observation. As can be seen from Fig. 8, as the value of $m_{\Delta}$ increases, the error decreases rapidly and reaches the machine precision after a certain value $\hat{m}_{\Delta}$, while the value of $\hat{m}_{\Delta}$ is independent of $M$ and $\omega$. In Fig. 9, we further show how the error changes with $m_{\Delta}$ for different values of $\gamma$ and $T_{\Delta}$. It can be seen that the value of $\hat{m}_{\Delta}$ decreases with the increase of $T_{\Delta}$. When $\gamma\leq 1$, the value of $\hat{m}_{\Delta}$ at the corresponding $\hat{T}_{1}$ is very close. When $\gamma>1$, the value of $\hat{m}_{\Delta}$ at the corresponding $\hat{T}_{1}$ increases rapidly with the increase of $\gamma$. From the perspective of use, we are more concerned about the $\hat{m}_{\Delta}$ value near $\hat{T}_{1}$. We tested this and the results are listed in Table 2.
@@
$\gamma$ $8$ $4$ $2$ $1$ $\frac{1}{2}$ $\frac{1}{4}$ $\frac{1}{8}$
@@
$\hat{m}_{\Delta}$ 1820 480 58 17 18 18 19
(a) $M=1000$, $\gamma=2$, $T_{\Delta}=6$.
(b) $\omega=20$, $\gamma=2$, $T_{\Delta}=6$.
(a) $M=2000$, $\gamma=8$, $\omega=20$.
(b) $M=1000$, $\gamma=4$, $\omega=20$.
(c) $M=500$, $\gamma=2$, $\omega=20$.
(d) $M=500$, $\gamma=1$, $\omega=20$.
(e) $M=500$, $\gamma=1/2$, $\omega=20$.
(f) $M=500$, $\gamma=1/4$, $\omega=20$.
Figure 9: The approximation error against $m_{\Delta}$ for various $\gamma$ and $T_{\Delta}$.
@@
Figure 10: The function and its approximation errors against $m_{\Delta}$ for $M=500$, $\gamma=1$, $T_{\Delta}=6$.
It is also worth noting that for functions with internal oscillations and smooth boundaries, an increase in the number of $m_{\Delta}$ will cause the value of $\hat{T}_{2}$ to decrease. At this time, if the value of $M$ is not large enough, the approximation result using a larger $m_{\Delta}$ may be worse. This can be seen from the approximation results for the function $f(t)=\cos\left(\frac{100}{1+25t^{2}}\right)$ in Figure 10.
4.3 Computational complexity and parameter determination
The computational complexity of this algorithm consists of three parts:
  • Calculation of $c_{\tau}$. The computational effort in this step is mainly concentrated on the singular value decomposition of the matrix $A$, which is related to $m_{\Delta}$ and $\gamma$, which can be estimated as:
    • If $\gamma<1$, the computational complexity is
      $\mathcal{O}\left(m_{\Delta}^{2}n_{\Delta}\right)=\mathcal{O}\left(\frac{1}{\gamma}m_{\Delta}^{3}\right)$ (22)
    • If $\gamma\geq 1$, the computational complexity is
      $\mathcal{O}\left(m_{\Delta}n^{2}_{\Delta}\right)=\mathcal{O}\left(\frac{1}{\gamma^{2}}m_{\Delta}^{3}\right)$ (23)
    $O\left(L_{\Delta}\log(L_{\Delta})\right)$ (24)
  • Calculation of Fourier coefficients of $f_{c}$. It can be achieved by FFT, the computational complexity is
    $O\left(M\log(M)\right).$ (25)
Combining the test results in the previous two subsections, we know that if we want to approximate the function with frequency $\omega$ with machine precision, the parameters $m_{\Delta}$ and $T_{\Delta}$ need to satisfy
$m_{\Delta}\geq\hat{m}_{\Delta},\quad T_{\Delta}\geq\hat{T}_{1}.$ (26)
Accordingly, combined with formula (21), the parameters $L_{\Delta}$ and $M$ need to satisfy
$L_{\Delta}\geq\hat{L}_{\Delta}=:2\times\hat{T}_{1}\times\hat{m}_{\Delta},\quad M\geq\hat{M}=:\hat{T}_{1}\gamma\omega.$ (27)
Referring to the results in Table 1 and Table 2, the value of $\hat{M}$ is very close for different $\gamma$ and the value of $\hat{L}_{\Delta}$ is relatively close when $\gamma=1$ and $2$, but the value of $\hat{L}_{\Delta}$ is much larger when $\gamma$ takes other values. Further combined with the value of $\hat{m}_{\Delta}$, it can be seen that the computational complexity of the method is equivalent when $\gamma=1$ and $\gamma=2$, while the computational complexity increases significantly in other cases. Furthermore, note that a larger $m_{\Delta}$ means a greater probability of encountering high frequencies in the selected boundary interval. Therefore, we recommend using the following parameters
$\gamma=1,T_{\Delta}=6,m_{\Delta}=20,$ (28)
@@ 4.3 Computational complexity and parameter determination
$\gamma=2,T_{\Delta}=2.3,m_{\Delta}=60,$ (29)
can be used as an alternative. Under such parameter settings, the size of the matrix $A$ in equation (15) is small, and we can directly calculate its singular value decomposition. In addition, note that $A$ is fixed in the algorithm, so the SVD of $A$ can be pre-computed and stored.
5 Testing and comparison of algorithm performance
Through the previous tests, we have given specific values for the parameters involved in the algorithm. In this section, we will further test the performance of the algorithm. All the tests are computed on Windows 10 system with Memory 16 GB, CPU Intel(R) Core(TM)i7-8500U CPU1.80GHz by MATLAB 2016b. We will compare the boundary interval algorithm (M1) proposed in this paper with Lyons algorithm (M2) and the Implicit projection algorithm (M3) in [14]. In the other two algorithms, we take the parameters as $T=2,\gamma=2$.
Figure 11 shows execution time for increasing the number of nodes $M$ of three algorithms. According to the parameter settings in (28), the algorithm only needs to perform an SVD of a matrix with a dimension of $40\times 41$, an IFFT of length $L_{\Delta}=2\times 6\times 20=240$, and an FFT of length $12M+$ 1. As $M$ increases, the speed of the present algorithm becomes apparent.
Figure 11: Execution time for increasing the number of nodes $M$, for MATLAB implementations of the Boundary Interval algorithm (M1), the Lyon algorithm (M2) and Implicit projection algorithm (M3).
The accuracy of the solution obtained through three algorithms is shown in Fig. 12 through 14. In order to better show the differences, we divide the function into three cases for comparison:
  • Overall smooth functions:
    $f_{1}(t)=t^{2},\quad f_{2}(t)=\exp(t),\quad f_{3}(t)=e^{\sin(2.7\pi t)+\cos(\pi t)}.$
  • Internal oscillating functions:
  • Boundary oscillating functions:
    $f_{7}(t)=\frac{1}{1.1-t^{2}},\quad f_{8}(t)=\text{Ai}(-66-70t),\quad f_{9}(t)=e^{\sin(65.5\pi t-27\pi)-\cos(20.6\pi t)}.$
(a) $f(t)=t^{2}$
@@ 5 Testing and comparison of algorithm performance
As can be seen from Figure 12, for the first case, when the function is overall smooth, the convergence of the present algorithm is slightly better than the other two algorithms, and it reaches the ideal accuracy at a smaller $M$. At this time, the number of nodes $M$ required by the three algorithms to achieve machine accuracy is relatively small, so there is no obvious difference in performance.
But as shown in Figure 13, for the second case, that is, when the function oscillates internally, since the method proposed in this article relies more on boundary data, the impact of this singularity on it is much smaller than the other two algorithms. The number of nodes $M$ required to achieve machine accuracy is significantly reduced (about $1/4$ of the other two algorithms).
For the third case, since the function oscillates at the boundary, the algorithm in this paper cannot avoid this singularity, so we see that the convergence performance of the three algorithms is very close. However, it should be noted that the number of nodes $M$ required by the three algorithms to achieve machine accuracy is very large at this time. Therefore, as shown in Figure 3, this method still has an advantage in computational complexity.
@@
(a) $f(t)=\frac{1}{1.1-t^{2}}$
@@ 6 Conclusions
A new fast algorithm based on boundary interval data has been proposed for the computation of the Fourier extension approximation to nonperiodic functions. We tested the relevant parameters and provided the setting scheme. The final algorithm has a computational complexity of $O(M\log(M))$ and can reduce the impact of internal singularities on reconstruction. This algorithm can be used as a basic tool for Fourier extension calculation.

### v4

@@ 5.2 Testing the parameter $m_{\Delta}$
The size of the parameter $m_{\Delta}$ determines whether this study is meaningful. This subsection tests the influence of the other parameters on it. First, we take $M>{T}_{\Delta}\gamma_{\Delta}\omega$ for observation. As shown in Fig. 8, as the value of $m_{\Delta}$ increases, the error decreases rapidly and reaches the machine precision after a certain value $\hat{m}_{\Delta}$, while the value of $\hat{m}_{\Delta}$ is independent of $M$ and $\omega$.
(a) $M=1000$, $\gamma_{\Delta}=2$, $T_{\Delta}=6$.
(b) $\omega=20$, $\gamma_{\Delta}=2$, $T_{\Delta}=6$.
(a) $M=2000$, $\gamma_{\Delta}=8$, $\omega=20$.
(b) $M=1000$, $\gamma_{\Delta}=4$, $\omega=20$.
(c) $M=500$, $\gamma_{\Delta}=2$, $\omega=20$.
(d) $M=500$, $\gamma_{\Delta}=1$, $\omega=20$.
(e) $M=500$, $\gamma_{\Delta}=1/2$, $\omega=20$.
(f) $M=500$, $\gamma_{\Delta}=1/4$, $\omega=20$.
Figure 9: The approximation error against $m_{\Delta}$ for various $\gamma_{\Delta}$ and $T_{\Delta}$.
In Fig. 9, we further show how the error changes with $m_{\Delta}$ for different values of $\gamma_{\Delta}$ and $T_{\Delta}$. It can be seen that the value of $\hat{m}_{\Delta}$ decreases with the increase of $T_{\Delta}$. This is consistent with the previous analysis, because the step size $h=\frac{2\pi}{L_{\Delta}}$ in the $g_{c}$ calculation process is determined by the product $(m_{\Delta}-1)T_{\Delta}$. When $\gamma_{\Delta}\leq 1$, the value of $\hat{m}_{\Delta}$ at the corresponding $\hat{T}_{1}$ is very close. When $\gamma_{\Delta}>1$, the value of $\hat{m}_{\Delta}$ at the corresponding $\hat{T}_{1}$ increases rapidly with the increase of $\gamma_{\Delta}$. From the perspective of use, we are more concerned about the $\hat{m}_{\Delta}$ value near $\hat{T}_{1}$. We tested this and the results are listed in Table 2.
@@
$\gamma_{\Delta}$ $8$ $4$ $2$ $1$ $\frac{1}{2}$ $\frac{1}{4}$ $\frac{1}{8}$
@@
$\hat{m}_{\Delta}$ 1870 490 64 23 25 24 26
@@
Figure 10: The function and its approximation errors against $m_{\Delta}$ for $M=500$, $\gamma_{\Delta}=1$, $T_{\Delta}=6$.
It is also worth noting that for functions with internal oscillations but are smooth near the boundary, as mentioned in Section 5, an increase in the number of $m_{\Delta}$ will cause the value of $\omega$ in (28) to increase. If the value of $M$ is not sufficiently large, the approximation result using a larger $m_{\Delta}$ may be worse. This can be seen from the approximation results for the function $f(t)=\cos\left(\frac{100}{1+25t^{2}}\right)$ in Fig. 10.
5.3 Parameter determination
Combining the test results in the previous two subsections, we know that if we want to approximate the function with frequency $\omega$ with machine precision, the parameters $T_{\Delta}$ and $m_{\Delta}$ must satisfy
$T_{\Delta}\geq\hat{T}_{1},\quad m_{\Delta}\geq\hat{m}_{\Delta}.$ (32)
Accordingly, combined with (31), the parameters $L_{\Delta}$ and $M$ must satisfy
$L_{\Delta}\geq\hat{L}_{\Delta}=:2\times\hat{T}_{1}\times\hat{m}_{\Delta},\quad M\geq\hat{M}=:\hat{T}_{1}\gamma_{\Delta}\omega.$ (33)
Referring to the results in Table 1 and Table 2, the value of $\hat{M}$ is very close for different $\gamma_{\Delta}$ values, and the value of $\hat{L}_{\Delta}$ is relatively close when $\gamma_{\Delta}=1$ and $2$; however, the value of $\hat{L}_{\Delta}$ is much larger when $\gamma_{\Delta}$ takes other values. Further combined with the value of $\hat{m}_{\Delta}$, it can be seen that the computational complexity of the method is equivalent when $\gamma_{\Delta}=1$ and $\gamma_{\Delta}=2$, whereas the computational complexity increases significantly in other cases. Furthermore, a larger $m_{\Delta}$ indicates a greater probability of encountering high frequencies within the selected boundary interval. Therefore, we recommend using the following parameters
$\gamma_{\Delta}=1,T_{\Delta}=6,m_{\Delta}=25,$ (34)
@@ 5.3 Parameter determination
$\gamma_{\Delta}=2,T_{\Delta}=2.3,m_{\Delta}=65,$ (35)
can be used as an alternative. Under such parameter settings, the size of matrix $A$ in (15) is small, and singular value decomposition can be calculated directly. In addition, note that $A$ is fixed in the algorithm, so the SVD of $A$ can be pre-computed and stored.
6 Testing and further improvement of the algorithm
6.1 Testing and comparison of algorithm performance
Figure 11: Execution time for increasing the number of nodes $M$, for MATLAB implementations of the Boundary Interval algorithm (M1), the FC-Gram algorithm (M2), the Lyon algorithm (M3) and Implicit projection algorithm (M4).
Based on the previous tests, we have provided specific values for the parameters involved in the algorithm. In this subsection, the performance of the proposed algorithm is tested. All tests were computed on a Windows 10 system with 16 GB of memory, Intel(R) Core(TM)i7-8500U CPU1.80GHz using MATLAB 2016b. We compared the boundary interval algorithm (M1) proposed in this paper with the FC-Gram algorithm (M2) in [5], the Lyons algorithm (M3) in [12] and the implicit projection algorithm (M4) in [14]. For convenience of numerical implementation, for M3, we set parameters $T=2,\gamma=2$. For M4, we selected the best result from the three parameter combinations of $T=6,\gamma=1$, $T=2,\gamma=2$, and $T=1.2,\gamma=4$ (the performance of the three parameter combinations for different functions is slightly different. We also tested other parameter combinations and found that no combination with a product of $T$ and $\gamma$ less than $4$ can achieve machine precision).
Algorithms M1 and M2 exhibit computational complexity of $\mathcal{O}(M\log M)$ , in contrast to the $\mathcal{O}(M\log^{2}M)$ complexity inherent to algorithms M3 and M4. Figure 11 demonstrates the scaling behavior of execution time versus node count $M$ across all four algorithms. These experimental observations align with the above conclusions: M1 and M2 demonstrate nearly identical temporal profiles, as do M3 and M4; for equivalent $M$ values, M1/M2 achieve superior computational efficiency compared with M3/M4. It should be noted that $M$ values required by these four algorithms are different. For M1/M2, the minimum $M$ is governed by the approximation capacity of the boundary interval and maximum frequency of the approximated function. For full-data algorithms such as M3 and M4, $M$ depends on both the function frequency and the critical parameter product $T\gamma$. Notably, our empirical studies and existing literature establish $T\gamma=4$ as the minimum allowable value for full-data algorithms to attain machine precision. Subthreshold $T\gamma$ values result in accuracy degradation for such methods.
(a) $f(t)=\text{erf}(2t)$
@@ 6.1 Testing and comparison of algorithm performance
The accuracy of the solution obtained using the four algorithms is shown in Fig. 12 - 14. To better show the differences, we divided the function into three cases for comparison:
  • Overall smooth functions:
    $f_{1}(t)=\text{erf}(2t),\quad f_{2}(t)=\text{Ai}(1+3t),\quad f_{3}(t)=e^{\sin(2.7\pi t)+\cos(\pi t)}.$
  • Internal oscillating functions:
  • Boundary oscillating functions:
    $f_{7}(t)=\cos(100t^{2}),\quad f_{8}(t)=\text{Ai}(-66-70t),\quad f_{9}(t)=e^{\sin(65.5\pi t-27\pi)-\cos(20.6\pi t)}.$
(a) $f(t)=\cos(100t^{2})$
@@
Fig. 12 shows the numerical results for the first case. As mentioned above, the number of nodes $M$ required for M1 and M2 is determined by the frequency of the boundary interval, but M2 cannot achieve superalgebraic convergence using polynomial approximation, so the value of $M$ required is significantly larger than M1; while M3 and M4 are determined by the maximum frequency of the function in the entire interval, and the number of nodes required is slightly larger than M1. At this time, the number of nodes $M$ required by the four algorithms to achieve machine accuracy was relatively small, therefore, there was no obvious difference in performance.
In the second case, that is, when the function oscillates internally. As the frequency of the function in the boundary interval is much lower than that in the interior, the number of nodes required for M1 and M2 is determined by the resolution constant of the standard Fourier transform. Note that the resolution constants of M3 and M4 are $T\gamma$, so from (29) we can see that the number of nodes required for M1and M2 is approximately $1/4$ of that for M3 and M4, which is consistent with the results given in Fig. 13.
In the third case, because the function oscillates near the boundary, the algorithm in this paper cannot avoid this singularity. For functions $f_{7}$ and $f_{8}$, the maximum frequency of the function appears in the boundary interval. According to (27), the number of nodes required by the proposed algorithm is $T_{\Delta}\gamma_{\Delta}=6$ times the maximum frequency, whereas the number of nodes required by M3 and M4 is $T\gamma=4$ times the maximum frequency. This is consistent with the results presented in Figures 2 and 3. For $f_{9}$, because the internal frequency is higher, the number of nodes required by the proposed algorithm is less than that of M3 and M4. As for M2, we can see that its approximation ability for the boundary interval data is obviously weaker than that of M1, therefore, more nodes are needed to achieve similar accuracy.
6.2 Algorithm improvement: boundary grid refinement
In certain situations, different step sizes may be used for interior and boundary sampling. The test results in the previous subsection indicate that the resolution constant is affected by the oscillation of the boundary interval data. This naturally raises the question: Can the algorithm performance be improved if a finer grid is used near the boundary than inside? In this section, we combine this concept to provide a modified algorithm for the boundary oscillation function.
Let $R$ be an positive integer and
${t}^{R}_{\ell}=\frac{\ell}{RM},\quad\ell=-RM,\ldots,RM.$ (36)
The positive integer $m^{R}_{\Delta}$ denotes the number of nodes selected in the boundary interval. Notably, only $2\times m^{R}_{\Delta}$ nodes near the boundary were used in the actual calculation. The other nodes can be ignored, but they are listed here for convenience. For consistency with the boundary interval taken in the previous algorithm, we let
${m}^{R}_{\Delta}=R(m_{\Delta}-1)+1$ (37)
and
$L_{\Delta}^{R}=2\times\left\lceil T_{\Delta}\times(m^{R}_{\Delta}-1)\right\rceil,h^{R}=\frac{2\pi}{L^{R}_{\Delta}},\quad x_{j}^{R}=(j-1)h,\quad j=1,2,\ldots,L^{R}_{\Delta}.$ (38)
Similarly, let
$J^{R}_{\Delta,1}=\{1,2,\ldots,m^{R}_{\Delta}\},~{}J^{R}_{\Delta,2}=\{\frac{L^{R}_{\Delta}}{2}+1,\frac{L^{R}_{\Delta}}{2}+2,\ldots,\frac{L^{R}_{\Delta}}{2}+m^{R}_{\Delta}\},~{}J^{R}_{\Delta}=J^{R}_{\Delta,1}\cup J^{R}_{\Delta,2}.$ (39)
and for $j\in J^{R}_{\Delta}$, take
$g^{R}(x^{R}_{j})=\left\{\begin{aligned} &f(t^{R}_{RM-m^{R}_{\Delta}+j}),&j\in J_{\Delta,1},\\ &f(t^{R}_{-RM+j-{L^{R}_{\Delta}}/{2}-1}),&j\in J_{\Delta,2}.\end{aligned}\right.$ (40)
The extension function $g^{R}_{c}(x^{R}_{j}),j=1,2,\ldots,L^{R}_{\Delta}$ was also obtained according to the method described in section 3. And then we obtain a function $f^{R}_{c}$ with a period of $2+\lambda$, where $\lambda$ is same as in (9). The data of the extension function $f^{R}_{c}(t_{\ell}),\ell=-M,\ldots,M,M+1,\ldots,M+\lceil T_{\Delta}-1\rceil\times(m_{\Delta}-1)-1$ in one period $[-1,1+\lambda)$ are
$f^{R}_{c}(t_{\ell})=\left\{\begin{aligned} &f(t_{\ell}),&|\ell|\leq M,\\ &g^{R}_{c}(x_{m^{R}_{\Delta}+(\ell-M)\cdot R}),&\ell>M.\end{aligned}\right.$ (41)
The original algorithm is equivalent to a special case of the improved algorithm when the parameter $R=1$.
Figure 15: The approximation error against $R$ for functions $\exp(\text{i}\omega\pi t)$, $M=500$.
(a) $f(t)=\frac{1}{1.01-t^{2}}$
(b) approximation error
(c) $f(t)=\text{Ai}(150t)$
(d) approximation error
(e) $f(t)=\sin(1500t^{2})$
(f) approximation error
Figure 16: Comparisons of algorithms M1 and M4.
Now we take $\gamma=1,T_{\Delta}=6,m_{\Delta}=25$ to test the influence of parameter $R$ on the algorithm. As can be seen from Fig. 15, when $R\geq 5$, the approximation error not only does not improve, but becomes worse, which is caused by the instability of the calculation of the extension part. When $R\leq 4$, it can be observed that the refinement of grid near the boundary improves the approximation error. When $\omega=50$, the improvement is not obvious, because the number of nodes $M=500>T_{\Delta}\gamma_{\Delta}\omega$ has exceeded the threshold for resolving this frequency function. Therefore, when facing high-frequency functions, we recommend using parameters
$\gamma_{\Delta}=1,T_{\Delta}=6,m_{\Delta}=25,R=4$
for calculation.
Next, we tested the performance of the improved algorithm (M5) and compared it with the original algorithm (M1). We selected several test functions with significant high-frequency features in the boundary interval:
$f_{10}(t)=\frac{1}{1.01-t^{2}},$ $f_{11}(t)=\text{Ai}(150t),$ $f_{12}(t)=\sin(1500t^{2}).$
Fig. 16 shows the graphes of the test functions and the corresponding approximation results. We can see that the refinement of grid near the boundary effectively reduces the resolution constant. The number of nodes required to achieve machine accuracy in the improved method was approximately $1/4$ of that in the original method.
@@ 7 Conclusions
Fast algorithms based on boundary interval data has been proposed for the computation of the Fourier extension approximation to nonperiodic functions. We tested the relevant parameters and provided the setting scheme. The new algorithms have a computational complexity of $\mathcal{O}(M\log(M))$ and can be used as basic tools for the calculation of the Fourier extension.
Acknowledgments
The authors would like to sincerely thank Mark Lyon for generously sharing the MATLAB codes in [5] and [12] to facilitate comparative analysis in this work.